what time will be required for a current of 0.1A to deposit all copper from 50dmcube of 0.1N solution? 

Grade:

1 Answers

bhanu p
12 years ago

we know that , N = W/E  => E = W/N  also , we have Z = E/96500  substituting this in faradays first law ,-

W = ZIt => W = EIt/96500 = WIt/N(96500)  => W = W(0.1)t/0.1(96500) => t = 96500(50) = 482500s